What monthly payment (to the nearest dollar) is required to pay off a $14000 car loan at 12% annual interest in 48 months?
Answer by josmiceli(6783)

What I come up with is
Annual Interest paid over 4 years (48 mos.) = Prt
P = principle
r = percentage rate
t = time (years)
I = 14000 * (.12) * 4
I = 6720 Interest paid in 4 years
If the interest payments are spread evenly over 4 years (48 mos.),
I(1 mo.) = 6720/48
I(1 mo.) = 140
The amount of the principle paid each month is
P(1 mo.) = 14000/48 = 291.67
I + P = $432 (to nearest dollar)
check
432 * (48) = 20736 (I + P paid in 48 mos.)
20736 - (I paid in 48 mos.) = P paid in 48 mos.
20736 - 6720 = 14016
(inaccuracy due to roundoff of answer)